write the slope-intercept equation for the line that passes through (10,-4) and has slope-3/10
Answer by Adrish(44)

Given: Slope=-3/10 and points (10,-4)
Slope intercept equation:
y= mx + b
substitute the value of slope and the value of x and y in the equation
-4=-3/10 . 10 +b
-4=-3+b
b=-4+3
b=-1
substitute the value of b in the equation
y= -3/10x -1